# Generic regularity of intermediate complex structure limits

### v1

@@ 1. Introduction
where the sections $F_{i}\in H^{0}(M,d_{i}L)$ and $F\in H^{0}(M,dL)$ define smooth and transverse divisors, and $d=\sum_{i=0}^{m}d_{i}$. Our goal is to improve the $C^{0}$-potential theoretic convergence to metric convergence. The non-Archimedean potential defines a family of ansatz metrics $\omega_{t}$ in the class $c_{1}(L)$, on some large open subset $U_{t}\subset X_{t}$ whose percentage of Calabi-Yau measure tends to one. We call $U_{t}$ the “generic region”, and we allow for $U_{t}$ to be slighted shrunken in the arguments.
@@
Acknowledgments. The first-named author was supported by a Royal Society University Research Fellowship. The second-named author was partially supported by NSF grant DMS-2404599.
@@ 2.2. Ricci-flat metrics
as $t\to 0$, see [2]. The Ricci-flat metrics $\omega_{{\rm CY},t}$ thus have volume form equal to
@@
where $\int_{X_{t}}\omega_{{\rm CY},t}^{n}=\int_{X_{t}}c_{1}(L)^{n}=:(L^{n})$ is independent of $t$. We will refer to this as the Calabi-Yau measure.
@@ 2.3. Ansatz reference metrics
Using the function $u$, [10, Corollary 4.6] defines open subsets $\Delta_{k}\subset\Delta,0\leqslant k\leqslant m,$ whose union has full measure in $\Delta$, and $u$ is smooth on each $\Delta_{k}$ by [10, Lemma 4.7]. Futhermore, [10, Lemma 4.8] shows that there is an open neighborhood $U_{k}\supset\Delta_{k}$ where the function $u$ does not depend on the coordinate $x_{k}$, and as a function on $\Delta_{k}$, the function $u$ satisfies the real Monge-Ampère PDE
@@
Variants of this PDE have also featured in recent works on complete noncompact Calabi-Yau manifolds [3, 4]. We emphasize that because $u$ does not depend on $x_{k}$ on $U_{k}$, it can be locally regarded as a function of $n$ variables. The Hessian $D^{2}u$ is taken with respect to these $n$ variables, and $D_{k}u=0$.
@@
where $d=\sum_{i=0}^{m}d_{i}$.
@@
is Calabi-Yau, and a neighborhood of $Y$ inside $X_{t}\subset M$ contains almost all the Calabi-Yau measure of $X_{t}$. Let $0\leqslant k\leqslant m$, and $K_{k}$ be compact subsets properly contained in the interior of $\Delta_{k}$, and we can assume its complement has arbitrarily small Lebesgue measure. We allow ourselves to slightly shrink $U_{k}$ to an open neighborhood of $K_{k}$, and via the logarithmic maps defined by $(x_{0},\ldots x_{k-1},x_{k+1},\ldots x_{m})$, this $U_{k}$ defines some open subset of $X_{t}$. Taking the union over $k=0,1,\ldots m$, we obtain the generic region $U_{t}\subset X_{t}$, whose normalized Calabi-Yau measure is arbitrarily close to one as $t$ approaches zero.
@@ 4. De Giorgi type blow up argument
then there is a quadratic polynomial $Q_{\eta r}$ on $\mathbb{R}^{m}$ of the form
@@
as long as the oscillations remain sufficiently small so that one can keep on iterating.
@@ 5. Higher order estimates on the generic region
Finally, we sketch how to modify this to prove the main Theorem 1.1. We recall that we are allowed to slightly shrink the “generic region” $U_{t}$ in the argument, as long as the normalized Calabi-Yau measure of its complement is smaller than any prescribed $\delta>0$. We can find compact subsets $K_{0},\ldots K_{m}$ properly contained inside the interior of $\Delta_{0},\ldots\Delta_{m}$, such that its complement has Lebesgue measure $\ll\delta$. Then $r_{0}=\text{dist}(K_{k},\partial\Delta_{k})>0$ is a small number independent of $t$, and we can define $U_{t}$ as above. By the main result of [10], the $C^{0}$-norm $\|\psi\|_{C^{0}(X_{t})}\to 0$ as $t\to 0$, so $\|\psi\|_{C^{0}(X_{t})}$ is far smaller than the constant $\varepsilon_{0}$, and we can run the De Giorgi type blow up argument, to prove Corollary 4.3. We can then follow the proof of Corollary 5.1, noting that the change of the Calabi-Yau volume form from the model to the geometric case only results in $O(e^{-cT})$-error, which is negligible. The conclusion is that up to slightly shrinking the compact sets $K_{0},\ldots K_{m}$, then for any $y\in K_{k}$ ($k=0,1,\ldots m$),

### v2

@@ 1. Introduction
where the sections $F_{i}\in H^{0}(M,d_{i}L)$ and $F\in H^{0}(M,dL)$ define smooth and transverse divisors, and $d=\sum_{i=0}^{m}d_{i}$. Our goal is to improve the $C^{0}$-potential theoretic convergence to metric convergence. The non-Archimedean potential defines a family of ansatz metrics $\omega_{t}$ in the class $c_{1}(L)$, on some large open subset $U_{t}\subset X_{t}$ whose percentage of Calabi-Yau measure tends to one. We call $U_{t}$ the “generic region”, and we allow for $U_{t}$ to be slightly shrunken in the arguments.
@@
Acknowledgments. We are grateful to the referee for useful comments. The first-named author was supported by a Royal Society University Research Fellowship. The second-named author was partially supported by NSF grant DMS-2404599.
@@ 2.2. Ricci-flat metrics
as $t\to 0$, see [2]. The Ricci-flat metrics $\omega_{{\rm CY},t}$ have normalized volume form equal to
@@
where $\int_{X_{t}}\omega_{{\rm CY},t}^{n}=\int_{X_{t}}c_{1}(L)^{n}=:(L^{n})$ is independent of $t$. We will refer to this as the normalized Calabi-Yau measure.
@@ 2.3. Ansatz reference metrics
Using the function $u$, [10, Corollary 4.6] defines open subsets $\Delta_{k}\subset\Delta,0\leqslant k\leqslant m,$ whose union has full measure in $\Delta$, and $u$ is smooth on each $\Delta_{k}$ by [10, Lemma 4.7]. Futhermore, [10, Lemma 4.8] shows that there are disjoint open neighborhoods $\Delta_{k}\subset U_{k}\subset\mathbb{R}^{m+1}$ where the function $u$ does not depend on the coordinate $x_{k}$, and as a function on $\Delta_{k}$, the function $u$ satisfies the real Monge-Ampère PDE
@@
(We note that $\Delta_{k}$ are open in $\Delta$, so their neighborhoods may be taken as disjoint). Variants of this PDE have also featured in recent works on complete noncompact Calabi-Yau manifolds [3, 4]. We emphasize that because $u$ does not depend on $x_{k}$ on $U_{k}$, it can be locally regarded as a function of $n$ variables. The Hessian $D^{2}u$ is taken with respect to these $n$ variables, and $D_{k}u=0$.
@@
where $d=\sum_{i=0}^{m}d_{i}$. In particular, on the region with uniformly bounded $\log|F|_{h^{d}}$, the image is $O(\frac{1}{|\log|t||})$-close to the standard $m$-simplex.
@@
is Calabi-Yau, and a neighborhood of $Y$ inside $X_{t}\subset M$ contains almost all the normalized Calabi-Yau measure of $X_{t}$. Let $0\leqslant k\leqslant m$, and $K_{k}$ be compact subsets properly contained in the interior of $\Delta_{k}$, and we can assume its complement has arbitrarily small Lebesgue measure. We allow ourselves to slightly shrink $U_{k}$ to an open neighborhood of $K_{k}$, and via the logarithmic maps defined by $(x_{0},\ldots x_{k-1},x_{k+1},\ldots x_{m})$, this $U_{k}$ defines some open subset of $X_{t}$. Taking the union over $k=0,1,\ldots m$, we obtain the generic region $U_{t}\subset X_{t}$, whose normalized Calabi-Yau measure is arbitrarily close to one as $t$ approaches zero.
@@ 4. De Giorgi type blow up argument
where $F$ is as in (3.16), then there is a quadratic polynomial $Q_{\eta r}$ on $\mathbb{R}^{m}$ of the form
@@
as long as $r/4^{k}\geqslant 100T^{-1/2}$ and the oscillations remain sufficiently small so that one can keep on iterating.
@@ 5. Higher order estimates on the generic region
Finally, we sketch how to modify this to prove the main Theorem 1.1. We recall that we are allowed to slightly shrink the “generic region” $U_{t}$ in the argument, as long as the normalized Calabi-Yau measure of its complement is smaller than any prescribed $\delta>0$. We can find compact subsets $K_{0},\ldots K_{m}$ properly contained inside the interior of $\Delta_{0},\ldots\Delta_{m}$, such that its complement has Lebesgue measure $\ll\delta$. Then $r_{0}=\text{dist}(K_{k},\partial\Delta_{k})>0$ is a small number independent of $t$, and we can define $U_{t}$ as above. By the main result of [10], the $C^{0}$-norm $\|\psi\|_{C^{0}(X_{t})}\to 0$ as $t\to 0$, so $\|\psi\|_{C^{0}(X_{t})}$ is far smaller than the constant $\varepsilon_{0}$, and we can run the De Giorgi type blow up argument, to prove Corollary 4.3. We can then follow the proof of Corollary 5.1, noting that the change of the normalized Calabi-Yau volume form from the model to the geometric case only results in $O(e^{-cT})$-error, which is negligible. The conclusion is that up to slightly shrinking the compact sets $K_{0},\ldots K_{m}$, then for any $y\in K_{k}$ ($k=0,1,\ldots m$),